A physics student drops a water balloon from a second floor window. The window is 18.05 m from the ground. How long does it take for the balloon to hit the ground?


1
Expert's answer
2020-10-30T13:19:40-0400

The time it will take is


"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2\cdot18.05}{9.8}}=1.919\text{ s}."